OpenStudy (anonymous):

Solve: 8x − 2(x + 1) = 7x + 8\ A.) 6 B.) -6 C.) 10 D.) -10

OpenStudy (anonymous):

x = -10

OpenStudy (anonymous):

8x-2x-2=7x+8 6x-2=7x+8 -2=x+8 D.) x=-10